Tags:
create new tag
view all tags

Shift Log 2012024

LN LC AA SB

17:00

Let the cavity Compton group continue after their half shift as they had some problems with their software earlier.

Plan: have Alex/Pavel optics, and some optics by Okugi-san to test for:

  • small beam size at LWIP (OTR measurement)
  • background levels in detector
  • response to skew tuning (OTR)

Can then use best optics for data taking in LW double shift later this week.

17:45 Start

Save settings: atf/data/set_data/set12apr24_1747.dat

Start trying Okugi-san's beta y* = 0.05 optics. Magnet changes:

  • for all of these sets: QM11FF: -37.804, QM12FF: -35.481, QM13FF: -10.207
  • beta y* = 0.05: QM14FF: -27.551, QM15FF: -36.238, QM16FF: 28.167, QF21X: 90.160.

Orbit tuning. Manipulator 6000, CV: 500, CH: 500.

18:20

Finding beam on OTR. Manipulator 6000, CH:500, CV: 500. Post IP OTR switched in.

Taking OTR shots of all optics sets:

beta y* QF21X QM16FF QM15FF QM14FF
0.01 88.736 27.722 -25.464 -59.179
0.02 90.591 28.302 -31.613 -45.495
0.03 90.704 28.337 -33.986 -37.592
0.04 90.486 28.269 -35.338 -31.960
0.05 90.160 28.167 -36.238 -27.551
0.06 89.792 28.052 -36.890 -23.902
0.07 89.404 27.931 -37.387 -20.767
0.08 89.007 27.807 -37.781 -18.001
0.09 88.603 27.681 -38.100 -15.510
0.10 88.195 27.553 -38.364 -13.227

Smallest σy ~ 2.2um at beta y* = 0.1, σx ~ 400um at this point.

20:00

Quad scan to see if this size and the large background (wire scanner detector saturated) can be reduced. QM14FF scanned. No improvement in beam size or background.

20:30

Returned to Alex/Pavel optics and tuning for background. After tuning background ~ 4000 counts with 10dB attenuation. Check beam size with OTR - decrease beam size with QM14FF. Apparent minimum at -78.73. Scan in 2.5A steps from -88.73 to -68.73. Fitted parabola minimum at -79.980A, σy = 2um.

Set file: setdata12apr24_2157.dat.

Polariser scan from -18 to 22 degrees. Min beam size σy = 1.59um, polariser angle 12.15 degrees.

23:00

Skew scan and log files.

QK4X log file
2.61 20120424_2253_log.dat
0 20120424_2256_log.dat
-5 20120424_2258_log.dat
-10 20120424_2300_log.dat
-15 20120424_2302_log.dat
-19 20120424_2304_log.dat
5 20120424_2306_log.dat
10 20120424_2308_log.dat
15 20120424_2310_log.dat
19 20120424_2311_log.dat

23:25

Delta f ramp

Delta f log file
+ 1kHz 20120424_2319_log.dat
+ 2kHz 20120424_2320_log.dat
+ 3kHz 20120424_2322_log.dat
+ 0kHz 20120424_2324_log.dat
- 1kHz 20120424_2325_log.dat
- 2kHz 20120424_2326_log.dat
- 3kHz 20120424_2327_log.dat

set file: set12apr24_2334.dat

23:50

Delta f ramp with OTR images

Background log (no ramp on): 20120424_2345_log.dat

Delta f log file
+ 1kHz 20120424_2350_log.dat
+ 2kHz 20120424_2352_log.dat
+ 3kHz 20120424_2354_log.dat
+ 0kHz 20120424_2348_log.dat
- 1kHz 20120424_2355_log.dat
- 2kHz 20120424_2357_log.dat
- 3kHz 20120424_2359_log.dat

500 pulse background log: 20120425_0003_log.dat

SB calibrated MFB2FF but may still be dodgy

  • QS1X and QS2X are both -0.595A

Delta f log file
-3kHz 20120424_0048_log.dat
-2kHz 20120424_0049_log.dat
-1kHz 20120424_0051_log.dat
0kHz 20120424_0052_log.dat
1kHz 20120424_0053_log.dat
2kHz 20120424_0054_log.dat
3kHz 20120424_0056_log.dat

  • QS1X and QS2X set to -0.495A

Delta f log file
-3kHz 20120424_0058_log.dat
-2kHz 20120424_0100_log.dat
-1kHz 20120424_0101_log.dat
0kHz 20120424_0102_log.dat
1kHz 20120424_0104_log.dat
2kHz 20120424_0105_log.dat
3kHz 20120424_0106_log.dat

  • QS1X and QS2X set to -0.395A

Delta f log file
-3kHz 20120424_0107_log.dat
-2kHz 20120424_0109_log.dat
-1kHz 20120424_0110_log.dat
0kHz 20120424_0111_log.dat
1kHz 20120424_0112_log.dat
2kHz 20120424_0113_log.dat
3kHz 20120424_0114_log.dat

  • QS1X and QS2X set to -0.695A

Delta f log file
-3kHz 20120424_0117_log.dat
-2kHz 20120424_0118_log.dat
-1kHz 20120424_0119_log.dat
0kHz 20120424_0120_log.dat
1kHz 20120424_0121_log.dat
2kHz 20120424_0122_log.dat
3kHz 20120424_0123_log.dat

* QS1X and QS2X set to -0.795A

Delta f log file
-3kHz 20120424_0125_log.dat
-2kHz 20120424_0126_log.dat
-1kHz 20120424_0127_log.dat
0kHz 20120424_0128_log.dat
1kHz 20120424_0130_log.dat
2kHz 20120424_0132_log.dat
3kHz 20120424_0134_log.dat

  • Magnet Movers at Start:
    20120424_1746_magmovers.jpg

  • Magnet Values at the start:
    20120424_1746_mags.jpg
Edit | Attach | Watch | Print version | History: r4 < r3 < r2 < r1 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r4 - 24 Apr 2012 - LaurieNevay

 
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback